OpenStudy (anonymous):

A nonzero product has 25 factors, 16 of which are negative. What is the sign of the products ? Explain

OpenStudy (ipwnbunnies):

16 negative factors. That means 25 - 16 = 9 factors are positive. Let's get rid of all the positive factors. If we multiplied each of the 9 positive factors by 9 negative factors, you should end up with a negative number. There are an odd amount of negative numbers being multiplied. We used up 18 factors, all 9 positive factors included. Now, we're left with 25 - 18 = 7 negative factors. We have our current negative numbers. If we multiplied 7 negative numbers, we'll get a positive number.

OpenStudy (anonymous):

So I can use that @iPwnBunnies

OpenStudy (ipwnbunnies):

Yeah, I suppose.

OpenStudy (anonymous):

Alright. @iPwnBunnies What number must be subtracted from -12 to obtain 24 ? 36 What number must be subtracted from -20 to obtain 15 ? 35 Is this right or wrong ?

OpenStudy (anonymous):

@iPwnBunnies

OpenStudy (ipwnbunnies):

It says subtracted from -12....hmm -12 - x = 24? That's how I see it in words. I'm not too sure though.

OpenStudy (anonymous):

So you don't know or you not sure @iPwnBunnies